### v1

# A Unique Bosonic Symmetry in a 4D Field-Theoretic System

Abstract: For the combined field-theoretic system of the four (3 + 1)-dimensional (4D) Abelian 3-form and 1-form gauge theories, we show the existence of a unique bosonic symmetry transformation that is constructed from the four infinitesimal, continuous and off-shell nilpotent symmetry transformations which exist for the Becchi-Rouet-Stora-Tyutin (BRST) quantized versions of the coupled (but equivalent) Lagrangian densities (that describe the present 4D field-theoretic system). The above off-shell nilpotent symmetry transformations are nothing but the BRST, co-BRST, anti-BRST and anti-co-BRST, under which, the Lagrangian densities transform to the total spacetime derivatives. The proof of the uniqueness of the above bosonic symmetry transformation operator crucially depends on the validity of all the four Curci-Ferrari (CF) type restrictions that exist on our theory. We highlight the importance of these CF-type restrictions, at various levels of our theoretical discussions, in the context of the unique bosonic transformation. We compare this observation against the backdrop of the three CF-type restrictions that appear in the requirements of the absolute anticommutativity of the specific set of nilpotent symmetries.
PACS numbers: 11.15.-q, 12.20.-m, 03.70.+k
Keywords: Combined field-theoretic system of the 4D free Ableian 3-form and 1-form gauge theories; off-shell nilpotent symmetries; a unique bosonic symmetry; CF-type restrictions
1 Introduction
It is an undeniable truth that the ideas of global and local symmetries (see, e.g. [1-3] and references therein) have played a key role in the development of theoretical physics (right from the very inception of physics as a branch of modern-day science). In particular, the 4D standard model of elementary particle physics, where there is a stunning degree of agreement between theory and experiment, owes a great deal to the interacting non-Abelian 1-from gauge theory which is based on the local Yang-Mills gauge symmetries. One of the key theoretical methods for the covariant canonical quantization of such kinds of theories (based on the gauge symmetries) is the Becchi-Rouet-Stora-Tyutin (BRST) formalism [5-7] where the unitarity and the quantum gauge (i.e. BRST) invariance are respected together at any arbitrary order of perturbative computations (see, e.g. [8] for details).
The central topic of our present endeavor is the four (3 + 1)-dimensional (4D) BRST-quantized version of the combined field-theoretic system of the free Abelian 3-form and 1-form gauge theories where we focus on the continuous as well as the discrete symmetry transformation operators and their algebraic structures. In particular, we concentrate on (i) the four continuous off-shell nilpotent symmetry transformation operators, (ii) a couple of discrete duality symmetry transformations, (iii) a unique continuous bosonic symmetry transformation operator (that is constructed from the appropriate anticommutator(s) between the above nilpotent symmetry transformation operators), and (iv) the derivation of the precise algebraic structures that are obeyed by the above continuous and discrete symmetry transformation operators. We demonstrate that there is an uncanny resemblance between the algebraic structures that are obeyed by (i) the symmetry transformation operators of our 4D BRST-quantized theory, and (ii) the de Rham cohomological operators of differential geometry (see, e.g. [9-11]). We also establish a two-to-one mapping between the symmetry transformation and cohomological operators***A set of three operators ($d,\;\delta,\;\Delta$), defined on a compact manifold without a boundary, is christened as the de Rham cohomological operators of differential geometry where $(\delta)d$ (with $\delta=\pm\,*\;d\;*,\;d=\partial_{\mu}\;(dx^{\mu}),\;\delta^{2}=0,\;d^{2}=0$) are the (co-)exterior derivatives and $\Delta=\{d,\;\delta\}$ stands for the Lalacian operator. Here the symbol * corresponds to the Hodge duality operator. These operators obey an algebra: $d^{2}=0,\;\delta^{2}=0,\;\Delta=\{d,\;\delta\},\;[\Delta,\;d]=0,\;[\Delta,\;\delta]=0$ which is popularly known as the Hodge algebra [9-11]. at the algebraic level.
Our present investigation is essential and important on the following counts. First of all, in our earlier works [12-14], we have been able to show the existence of a set of four infinitesimal, continuous and off-shell nilpotent (anti-)BRST and (anti-)co-BRST symmetry transformation operators for our present 4D BRST-quantized field-theoretic combined system of the free Abelian 3-form and 1-from gauge theories. We are curious to know the algebraic structure that is obeyed by these fermionic symmetry operators in our present endeavor. Second, we have also been able to demonstrate that our 4D theory respects a couple of discrete duality symmetry transformation operators. Thus, we are very eager to know their roles in the whole gamut of the algebraic structure that is respected by various kinds of symmetry transformation operators of our theory. Third, we lay emphasis on the roles of the CF-type restrictions on our 4D BRST-quantized theory (as far as the whole beauty of the algebraic structure is concerned). Fourth, out of the six anticommutators that exist amongst the four nilpotent symmetry transformation operators, it is very interesting to show that (i) the two of them absolutely anticmmute with each-other (modulo the U(1) vector gauge symmetry-type transformations), (ii) the other two anticommutators are found to be zero provided we use the specific set of three (out of the total four) CF-type restrictions that exist on our 4D theory, and (iii) the rest of the two anticomutators define the bosonic symmetry transformation operators. Finally, we establish that, out of the above two bosonic symmetry transformation operators, only one of them is independent on the space of quantum fields where all the four CF-type restrictions are satisfied.
The key theoretical contents of our present endeavor are organized as follows. In the next section, we recapitulate the bare essentials of the four nilpotent transformations that exist for our 4D BRST-quantized field-theoretic system [13,14]. In our section three, we discuss the algebraic structure of the discrete and continuous symmetry transformation operators and define a set of two bosonic symmetry transformation operators from the four nilpotent symmetry transformation operators. In particular, we point out the appearance of a specific set of three (out of the total four) CF-type restrictions in the context of the requirement of the absolute anticommutativity between (i) the (anti-)BRST symmetry transformation operators, and (ii) the (anti-)co-BRST symmetry transformation operators. Our section four deals with the bosonic symmetry transformations on (i) the basic fields of our theory, and (ii) the coupled Lagrangian densities, where we show (i) the appearance of the full set of the four CF-type restrictions, and (ii) the proof of the uniqueness of the bosonic symmetry transformation operator. Finally, in our section five, we summarize our key observations and point out the future perspective of our present investigation.
2 Preliminaries: Off-Shell Nilpotent Symmetries
[…]
- $\,\dfrac{1}{4}\,\big(B_{\mu\nu}\big)^{2}+\dfrac{1}{2}\,B_{\mu\nu}\,\Big[\partial_{\sigma}A^{\sigma\mu\nu}+\dfrac{1}{2}\,\big(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu}\big)\Big]-\dfrac{1}{4}\,B_{2}^{2}+\dfrac{1}{2}\,B_{2}\big(\partial\cdot\phi\big)$
$+$ $\dfrac{1}{4}\,\big({\cal B}_{\mu\nu}\big)^{2}-\dfrac{1}{2}\,{\cal B}_{\mu\nu}\,\Big[\varepsilon^{\mu\nu\sigma\rho}\,\partial_{\sigma}A_{\rho}+\dfrac{1}{2}\,\big(\partial^{\mu}\widetilde{\phi}^{\nu}-\partial^{\nu}\widetilde{\phi}^{\mu}\big)\Big],$
${\cal L}_{(FP)}$ $=$ $\dfrac{1}{2}\,\Big[\big(\partial_{\mu}\bar{C}_{\nu\sigma}+\partial_{\nu}\bar{C}_{\sigma\mu}+\partial_{\sigma}\bar{C}_{\mu\nu}\big)\big(\partial^{\mu}C^{\nu\sigma}\big)+\big(\partial_{\mu}\bar{C}^{\mu\nu}+\partial^{\nu}\bar{C}_{1}\big)f_{\nu}$ (2)
- $\big(\partial_{\mu}C^{\mu\nu}+\partial^{\nu}C_{1}\big)\bar{F}_{\nu}+\big(\partial\cdot\bar{\beta}\big)\,B_{4}-\big(\partial\cdot\beta\big)\,B_{5}-B_{4}\,B_{5}-2\,\bar{F}^{\mu}\,f_{\mu}$
- $\big(\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\mu}\bar{\beta}_{\nu}\big)\big(\partial^{\mu}\beta^{\nu}\big)-\,\partial_{\mu}\bar{C}_{2}\,\partial^{\mu}C_{2}\Big]-\partial_{\mu}\bar{C}\,\partial^{\mu}C,$
where the Abelian 1-form ($A^{(1)}=A_{\mu}\,dx^{\mu}$) and 3-form $A^{(3)}=\frac{1}{3!}\,A_{\mu\nu\sigma}\,(d\,x^{\mu}\wedge d\,x^{\nu}\wedge d\,x^{\sigma})$] define the free basic gauge fieldsOur 4D flat spacetime manifold is endowed with the metric tensor $\eta_{\mu\nu}$ = diag (+1, -1, -1, -1) so that the dot product between two non-null vectors $P_{\mu}$ and $Q_{\mu}$ is: $(P\cdot Q)=\eta_{\mu\nu}P^{\mu}Q^{\nu}=P_{0}Q_{0}-P_{i}Q_{i}$ where the Greek indices $\mu,\nu,\sigma...=0,1,2,3$ and the Latin indeces i,j,k=1,2,3i,j,k...=1,2,3. We choose the 4D Levi-Civita tensor $\varepsilon_{\mu\nu\sigma\rho}$ such that $\varepsilon_{0123}=+1=-\varepsilon^{0123}$ and, when two of them are contracted together, the standard relationships: $\varepsilon_{\mu\nu\sigma\rho}\varepsilon^{\mu\nu\sigma\rho}=-\,4!,\;\varepsilon_{\mu\nu\sigma\rho}\varepsilon^{\mu\nu\sigma\eta}=-\,3!\,\delta^{\eta}_{\rho},\;\varepsilon_{\mu\nu\sigma\rho}\varepsilon^{\mu\nu\eta\kappa}=-\,2!(\delta^{\eta}_{\sigma}\delta^{\kappa}_{\rho}-\delta^{\eta}_{\rho}\delta^{\kappa}_{\sigma})$, etc., are satisfied. $A_{\mu}$ and $A_{\mu\nu\sigma}$, respectively. In the Lagrangian density ${\cal L}_{(NG)}$, the set of bosonic Nakanishi-Lautrup type auxiliary fields $B,B_{1},B_{2},B_{3},B_{\mu\nu},{\cal B}_{\mu\nu}$ have been invoked for the linearization purposes. For instance, the auxiliary fields $B_{1},\,{\cal B}_{\mu\nu}$ have been utilized in the linearizations of the kinetic terms for the Abelian 3-form and 1-form gauge fields, respectively. On the other hand, the set of auxiliary fields $B,(B_{3})B_{2},B_{\mu\nu}$ have been used to linearize the gauge-fixing terms for the Abelian gauge field $A_{\mu}$, the (axial-)vector fields $(\widetilde{\phi}_{\mu})\phi_{\mu}$ (which appear in the theory because of the reducibility properties) and the Abelian 3-form gauge field $A_{\mu\nu\sigma}$, respectively. In the Faddeev-Popov (FP) ghost part of the Lagrangian density ${\cal L}_{(FP)}$, we have the basic (anti-)ghost fields: $(\bar{C}_{\mu\nu})C_{\mu\nu},\,(\bar{\beta}_{\mu})\beta_{\mu},\,(\bar{C}_{2})C_{2},(\bar{C}_{1})C_{1},\,(\bar{C})C$ with the ghost numbers: (1)+1,(2)+2,(3)+3,(1)+1,(1)+1(-1)+1,(-2)+2,(-3)+3,$(-1)+1,$ respectively. The ferminonic (anti-)ghost fields $(\bar{C})C$ are associated with the Abelian 1-form gauge field and the rest of the above basic (anti-)ghost fields are connected to the Abelian 3-from gauge field $A_{\mu\nu\sigma}$. In the above, all the basic (anti-)ghost fields are fermionic in nature except the (anti-)ghost vector fields $(\bar{\beta}_{\mu})\beta_{\mu}$ which are bosonic (i.e. $\beta_{\mu}^{2}\neq 0,\;\bar{\beta}_{\mu}^{2}\neq 0,\,\beta_{\mu}\bar{\beta}_{\nu}=\bar{\beta}_{\nu}\beta_{\mu}$, etc.) in nature. We also have fermionic auxiliary (anti-)ghost fields $(\bar{F}_{\mu})f_{\mu}$ with the ghost numbers (-1)+1 and the bosonic set of auxiliary (anti-)ghost fields $(B_{5})B_{4}$ with the ghost numbers (-2)+2, respectively. It can be checked that the following infinitesimal, continuous and off-shell nilpotent (i.e. $s_{(d)b}^{2}=0$) (co-)BRST [i.e. (dual-)BRST] symmetry transformations $s_{(d)b}$, namely;
$s_{d}A_{\mu\nu\sigma}=\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\rho}\bar{C},\quad s_{d}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}\bar{C}^{\sigma\rho},\quad s_{d}\bar{C}_{\mu\nu}=\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\nu}\bar{\beta}_{\mu},\quad s_{d}\bar{f}_{\mu}=\partial_{\mu}B_{5},$
$s_{d}\bar{\cal B}_{\mu\nu}=\partial_{\mu}\bar{F}_{\nu}-\partial_{\nu}\bar{F}_{\mu},\quad s_{d}\bar{\beta}_{\mu}=\partial_{\mu}\bar{C}_{2},\quad s_{d}C_{1}=+\,B_{3},\quad s_{d}\beta_{\mu}=-\,f_{\mu},\quad s_{d}\widetilde{\phi}_{\mu}=+\,\bar{F}_{\mu},$
$s_{d}C_{\mu\nu}=+\,{\cal B}_{\mu\nu},\qquad s_{d}C=+\,B_{1},\quad s_{d}C_{2}=B_{4},\qquad s_{d}\bar{C}_{1}=B_{5},\qquad s_{d}F_{\mu}=+\,\partial_{\mu}B_{3},$
$s_{d}\;\Bigl[\bar{C}_{2},\,\bar{C},\,f_{\mu},\,{\bar{F}}_{\mu},\,\phi_{\mu},\,B,\,B_{1},\,B_{2},\,B_{3},\,B_{4},\,B_{5},\,B_{\mu\nu},\,\bar{B}_{\mu\nu},\,{\cal B}_{\mu\nu}\Bigl]\;=\;0,$ (3)
$s_{b}A_{\mu\nu\sigma}=\partial_{\mu}C_{\nu\sigma}+\partial_{\nu}C_{\sigma\mu}+\partial_{\sigma}C_{\mu\nu},\qquad s_{b}C_{\mu\nu}=\partial_{\mu}\beta_{\nu}-\partial_{\nu}\beta_{\mu},\qquad s_{b}\bar{C}_{\mu\nu}=B_{\mu\nu},$
$s_{b}A_{\mu}=\partial_{\mu}C,\quad s_{b}\bar{C}=B,\quad s_{b}\bar{\beta}_{\mu}=\bar{F}_{\mu},\quad s_{b}\beta_{\mu}=\partial_{\mu}C_{2},\quad s_{b}\bar{B}_{\mu\nu}=\partial_{\mu}f_{\nu}-\partial_{\nu}f_{\mu},$
$s_{b}{\bar{C}}_{2}=B_{5},\qquad s_{b}C_{1}=-B_{4},\qquad s_{b}\bar{C}_{1}=B_{2},\qquad s_{b}\phi_{\mu}=f_{\mu},\qquad s_{b}F_{\mu}=-\,\partial_{\mu}B_{4},$
$s_{b}\bar{f}_{\mu}=\partial_{\mu}B_{2},\;s_{b}\Bigl[C,\,C_{2},\,B,\,B_{1},\,B_{2},\,B_{3},\,B_{4},\,B_{5},\,f_{\mu},\,{\bar{F}}_{\mu},\,\widetilde{\phi}_{\mu},B_{\mu\nu},\,{\cal B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu}\Bigl]=0,$ (4)
leave the action integral $S=\int d^{4}x\,{\cal L}_{(B)}$, corresponding to the total Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$, invariant (i.e. $s_{(d)b}S=0$) because we observe the following:
$s_{d}{\cal L}_{(B)}$ $=$ $-\,\dfrac{1}{2}\,\partial_{\mu}\,\Big[(\partial^{\mu}\,\bar{C}^{\nu\sigma}+\partial^{\nu}\,\bar{C}^{\sigma\mu}+\partial^{\sigma}\,\bar{C}^{\mu\nu})\,{\cal B}_{\nu\sigma}+{\cal B}^{\mu\nu}\,\bar{F}_{\nu}-B_{4}\,\partial^{\mu}\,\bar{C}_{2}$ (5)
$+$ $B_{3}\,\bar{F}^{\mu}-B_{5}\,f^{\mu}-(\partial^{\mu}\,\bar{\beta}^{\nu}-\partial^{\nu}\,\bar{\beta}^{\mu})\,f_{\nu}\Big]+\partial_{\mu}\Big[B_{1}\,\partial^{\mu}\bar{C}\Big],$
$s_{b}\,{\cal L}_{(B)}$ $=$ $\dfrac{1}{2}\,\partial_{\mu}\,\Big[(\partial^{\mu}\,C^{\nu\sigma}+\partial^{\nu}\,C^{\sigma\mu}+\partial^{\sigma}\,C^{\mu\nu})\,B_{\nu\sigma}+B^{\mu\nu}\,f_{\nu}-B_{5}\,\partial^{\mu}\,C_{2}$ (6)
$+$ $B_{2}\,f^{\mu}+B_{4}\,\bar{F}^{\mu}-(\partial^{\mu}\,\beta^{\nu}-\partial^{\nu}\,\beta^{\mu})\,\bar{F}_{\nu}\Big]-\partial_{\mu}\Big[B\,\partial^{\mu}C\Big].$
Corresponding to the Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$ [cf. Eqs. (1),(2)], we have a coupled Lagrangian density ${\cal L}_{(\bar{B})}={\cal L}_{(ng)}+{\cal L}_{(fp)}$ whose non-ghost sector [i.e. ${\cal L}_{(ng)}$] and FP-ghost sector [i.e. ${\cal L}_{(fp)}$] parts (of the Lagrangian density) are
${\cal L}_{(ng)}=\dfrac{1}{2}\,B^{2}-B\,(\partial\cdot A)+\dfrac{1}{2}\,B_{2}\,\big(\partial\cdot\phi\big)-\dfrac{1}{4}\,B_{2}^{2}-\dfrac{1}{2}\,B_{3}\,\big(\partial\cdot\widetilde{\phi}\big)+\dfrac{1}{4}\,B_{3}^{2}$ (7)
$+$ $\dfrac{1}{2}\,B_{1}^{2}+\,B_{1}\,\Big(\dfrac{1}{3!}\,\varepsilon^{\mu\nu\sigma\rho}\,\partial_{\mu}A_{\nu\sigma\rho}\Big)-\,\dfrac{1}{4}\,\big(\bar{B}_{\mu\nu}\big)^{2}-\dfrac{1}{2}\,\bar{B}_{\mu\nu}\,\Big[\partial_{\sigma}A^{\sigma\mu\nu}+\dfrac{1}{2}\,\big(\partial^{\mu}\phi^{\nu}-\partial^{\nu}\phi^{\mu}\big)\Big]$
$+$ $\dfrac{1}{4}\,\big({\bar{\cal B}_{\mu\nu}}\big)^{2}+\dfrac{1}{2}\,{\bar{\cal B}_{\mu\nu}}\,\Big[\varepsilon^{\mu\nu\sigma\rho}\,\partial_{\sigma}A_{\rho}-\dfrac{1}{2}\,\big(\partial^{\mu}\widetilde{\phi}^{\nu}-\partial^{\nu}\widetilde{\phi}^{\mu}\big)\Big],$
${\cal L}_{(fp)}$ $=$ $\dfrac{1}{2}\,\Big[\big(\partial_{\mu}\bar{C}_{\nu\sigma}+\partial_{\nu}\bar{C}_{\sigma\mu}+\partial_{\sigma}\bar{C}_{\mu\nu}\big)\big(\partial^{\mu}C^{\nu\sigma}\big)-\big(\partial_{\mu}\bar{C}^{\mu\nu}-\partial^{\nu}\bar{C}_{1}\big)F_{\nu}$ (8)
$+$ $\big(\partial_{\mu}C^{\mu\nu}-\partial^{\nu}C_{1}\big)\bar{f}_{\nu}+\big(\partial\cdot\bar{\beta}\big)\,B_{4}-\big(\partial\cdot\beta\big)\,B_{5}-B_{4}\,B_{5}-2\,\bar{f}^{\mu}\,F_{\mu}$
- $\big(\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\mu}\bar{\beta}_{\nu}\big)\big(\partial^{\mu}\beta^{\nu}\big)-\,\partial_{\mu}\bar{C}_{2}\,\partial^{\mu}C_{2}\Big]-\partial_{\mu}\bar{C}\,\partial^{\mu}C,$
where a couple of new (i) bosonic auxiliary fields $(\bar{B}_{\mu\nu})\bar{\cal B}_{\mu\nu}$, and (ii) fermionic auxiliary fields $(F_{\mu})\bar{f}_{\mu}$ have been invoked in the Lagrangian densities ${\cal L}_{(ng)}$ and ${\cal L}_{(fp)}$, respectively. It is straightforward to check that the following off-shell nilpotent (i.e. $s_{ad}^{2}=0,\,s_{ab}^{2}=0$) anti-dual-BRST ($s_{ad}$) and anti-BRST ($s_{ab}$) symmetry transformations
$s_{ad}A_{\mu\nu\sigma}=\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\rho}C,\qquad s_{ad}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}C^{\sigma\rho},\qquad s_{ad}C_{\mu\nu}=-\,\big(\partial_{\mu}\beta_{\nu}-\partial_{\nu}\beta_{\mu}\big),$
$s_{ad}{\cal B}_{\mu\nu}=\partial_{\mu}F_{\nu}-\partial_{\nu}F_{\mu},\qquad s_{ad}\beta_{\mu}=-\,\partial_{\mu}C_{2},\qquad s_{ad}\bar{\beta}_{\mu}=+\,\bar{f}_{\mu},\quad s_{ad}\widetilde{\phi}_{\mu}=+\,F_{\mu},$
$s_{ad}\bar{C}_{\mu\nu}=+\,\bar{\cal B}_{\mu\nu},\qquad s_{ad}\bar{C}=-\,B_{1},\qquad s_{ad}\bar{C}_{2}=-\,B_{5},\qquad s_{ad}\bar{C}_{1}=-\,B_{3},$
$s_{ad}C_{1}=-\,B_{4},\qquad\quad s_{ad}f_{\mu}=-\,\partial_{\mu}B_{4},\qquad s_{ad}\bar{F}_{\mu}=-\,\partial_{\mu}B_{3},$
$s_{ad}\;\Bigl[C,\,C_{2},\,B,\,B_{1},\,B_{2},\,B_{3},\,B_{4},\,B_{5},\,\bar{f}_{\mu},\,{F}_{\mu},\,\phi_{\mu},\,B_{\mu\nu},\,\bar{B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu}\Bigl]\;=\;0,$ (9)
$s_{ab}A_{\mu\nu\sigma}=\partial_{\mu}\bar{C}_{\nu\sigma}+\partial_{\nu}\bar{C}_{\sigma\mu}+\partial_{\sigma}\bar{C}_{\mu\nu},\quad s_{ab}\bar{C}_{\mu\nu}=\partial_{\mu}\bar{\beta}_{\nu}-\partial_{\nu}\bar{\beta}_{\mu},\quad s_{ab}C_{\mu\nu}=\bar{B}_{\mu\nu},$
$s_{ab}A_{\mu}=\partial_{\mu}\bar{C},\quad s_{ab}C=-\,B,\quad s_{ab}\beta_{\mu}=F_{\mu},\quad s_{ab}\bar{\beta}_{\mu}=\partial_{\mu}\bar{C}_{2},\quad s_{ab}B_{\mu\nu}=\partial_{\mu}\bar{f}_{\nu}-\partial_{\nu}\bar{f}_{\mu},$
$s_{ab}{C}_{2}=B_{4},\quad s_{ab}C_{1}=-B_{2},\quad s_{ab}\bar{C}_{1}=-\,B_{5},\quad s_{ab}\phi_{\mu}=\bar{f}_{\mu},\quad s_{ab}\bar{F}_{\mu}=-\,\partial_{\mu}B_{5},$
$s_{ab}f_{\mu}=-\,\partial_{\mu}B_{2},\;s_{ab}\Bigl[\bar{C},\,\bar{C}_{2},\,B,\,B_{1},B_{2},B_{3},B_{4},B_{5},\bar{f}_{\mu},\,{F}_{\mu},\,\widetilde{\phi}_{\mu},\,\bar{B}_{\mu\nu},\,{\cal B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu}\Bigl]=0,$ (10)
render the action integral $S=\int d^{4}x\,{\cal L}_{\bar{B}}$ invariant (i.e. $s_{ad}S=0,\,s_{ad}S=0$) because of the following transformations of ${\cal L}_{(\bar{B})}$ under $s_{ad}$ and $s_{ab}$, namely; :
$s_{ad}\,{\cal L}_{(\bar{B})}$ $=$ $\dfrac{1}{2}\,\partial_{\mu}\,\Big[(\partial^{\mu}\,C^{\nu\sigma}+\partial^{\nu}\,C^{\sigma\mu}+\partial^{\sigma}\,C^{\mu\nu})\,\bar{\cal B}_{\nu\sigma}-\bar{\cal B}^{\mu\nu}\,F_{\nu}+B_{5}\,\partial^{\mu}\,C_{2}$ (11)
- $B_{3}\,F^{\mu}+B_{4}\,\bar{f}^{\mu}-(\partial^{\mu}\,\beta^{\nu}-\partial^{\nu}\,\beta^{\mu})\,\bar{f}_{\nu}\Big]+\partial_{\mu}\Big[B_{1}\,\partial^{\mu}C\Big],$
$s_{ab}\,{\cal L}_{(\bar{B})}$ $=$ $\dfrac{1}{2}\,\partial_{\mu}\,\Big[\bar{B}^{\mu\nu}\,\bar{f}_{\nu}-\,(\partial^{\mu}\,\bar{C}^{\nu\sigma}+\partial^{\nu}\,\bar{C}^{\sigma\mu}+\partial^{\sigma}\,\bar{C}^{\mu\nu})\,\bar{B}_{\nu\sigma}+B_{4}\,\partial^{\mu}\,\bar{C}_{2}$ (12)
- $B_{5}\,F^{\mu}+B_{2}\,\bar{f}^{\mu}-(\partial^{\mu}\,\bar{\beta}^{\nu}-\partial^{\nu}\,\bar{\beta}^{\mu})\,F_{\nu}\Big]-\partial_{\mu}\Big[B\,\partial^{\mu}\bar{C}\Big].$
Thus, the action integral of our 4D BRST-quantized field-theoretic system respects total four off-shell nilpotent symmetry transformations [cf. Eqs. (3),(4),(9),(10)].
We end our present section with the remarks that (i) the total kinetic terms for the Abelian 3-form and 1-form gauge fields remain invariant under the nilpotent (anti-)BRST symmetry transformations, (ii) the total gauge-fixing terms, corresponding to the Abelian 3-form and 1-form gauge fields, remain unchanged under the nilpotent (anti-)dual-BRST transformations, and (iii) the BRST-quantized versions of the Lagrangian densities [that exist for the free Abelian 3-form gauge theory in equations (1), (2), (7) and (8)] have been taken from our earlier work where they have been systematically derived (see, e.g. [15]).
3 Algebraic Structures: Discrete Duality and Continuous Nilpotent Symmetry Operators
In our present section, we focus on (i) the the discrete duality symmetry transformation operator, (ii) the continuous and nilpotent symmetry transformation operators, and (iii) the algebraic structures that are satisfied by them. In particular, we attach a great of importance to the emergence of the (anti-)BRST as well as the (anti-)co-BRST invariant CF-type restrictions in our present discussions. In this context, first of all, we note that the following discrete duality symmetry transformations for the basic and auxiliary fields§§§It is worthwhile to mention that our current discrete duality symmetry transformations (13) are slightly different from such transformations quoted in our earlier work [12] and these are the correct set because they satisfy accurately the algebraic structure that have been mentioned in (14) and (17).
$A_{\mu}\longrightarrow\,\mp\,\dfrac{1}{3!}\,\varepsilon_{\mu\nu\sigma\rho}\,A^{\nu\sigma\rho},\quad A_{\mu\nu\sigma}\longrightarrow\,\pm\,\varepsilon_{\mu\nu\sigma\rho}\,A^{\rho},\quad B_{\mu\nu}\to\pm\,{\cal B}_{\mu\nu},\quad{\cal B}_{\mu\nu}\to\mp\,B_{\mu\nu},$
$B\to\pm\,B_{1},\quad B_{1}\to\mp\,B,\quad B_{2}\to\pm\,B_{3},\quad B_{3}\to\mp\,B_{2},\;\;\phi_{\mu}\to\pm\,\widetilde{\phi}_{\mu},\;\;\widetilde{\phi}_{\mu}\to\mp\,\phi_{\mu},$
$C_{\mu\nu}\longrightarrow\,\pm\,\bar{C}_{\mu\nu},\quad\bar{C}_{\mu\nu}\longrightarrow\,\mp\,C_{\mu\nu},\quad\beta_{\mu}\to\mp\,\bar{\beta}_{\mu},\quad\bar{\beta}_{\mu}\to\mp\,\beta_{\mu},\quad f_{\mu}\to\pm\,\bar{F}_{\mu},$
$\bar{F}_{\mu}\to\mp\,f_{\mu},\qquad B_{4}\to\pm\,B_{5},\qquad B_{5}\to\pm\,B_{4},\qquad C\to\mp\,\bar{C},\qquad\bar{C}\to\pm\,C,$
$C_{2}\to\pm\,\bar{C}_{2},\qquad\bar{C}_{2}\to\mp\,C_{2},\qquad C_{1}\to\pm\,\bar{C}_{1},\qquad\bar{C}_{1}\to\mp\,C_{1},$ (13)
leave the Lagrangian density ${\cal L}_{(B)}={\cal L}_{(NG)}+{\cal L}_{(FP)}$ [cf. Eqs. (1),(2)] invariant. It is worthwhile to mention, at this stage, that the interplay between the continuous symmetry operators [cf. Eq. (3),(4)] and discrete symmetry operator [cf. Eq. (13)], provide the physical realization of the well-known operator relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives $(\delta)d$ of differential geometry because we observe the following
$s_{d}\,\Sigma$ $=$ $+\,*\,s_{b}\,*\,\Sigma,\qquad\Sigma=\beta_{\mu},\;\bar{\beta}_{\mu},\,B_{4},\,B_{5},\qquad s_{d}\,\Phi=-\,*\,s_{b}\,*\,\Phi,$
$\Phi$ $=$ $A_{\mu\nu\sigma},\,B_{\mu\nu},\,{\cal B}_{\mu\nu},\,\bar{C}_{\mu\nu},\,C_{\mu\nu},\,A_{\mu},\phi_{\mu},\,\widetilde{\phi}_{\mu},\,f_{\mu},\,\bar{F}_{\mu},$ (14)
$\bar{C},\,C,\,\bar{C}_{1},\,C_{1},\,\bar{C}_{2},\,C_{2},\,B,\,B_{1},\,B_{2},\,B_{3},$
where the $\pm$ signs in the relationships $s_{d}\,\Sigma=+\,*\,s_{b}\,*\,\Sigma$ and $s_{d}\,\Phi=-\,*\,s_{b}\,*\,\Phi$ are dictated by a couple of successive operations of the discrete duality symmetry transformations (13) on the generic fields $\Sigma$ and $\Phi$ which turn out to be the following (see, e.g. [16]):
$*\;\big(*\;\Sigma\big)=+\,\Sigma,\;\;\qquad\;\;*\;\big(*\;\Phi\big)=-\,\Phi.$ (15)
It is worthwhile to point out that the nilpotent (i.e. $\delta^{2}=0,\,d^{2}=0$) (co-)exterior derivative operators $(\delta)d$ are (i) physically realized [cf. Eq. (14)] in terms of the nilpotent (i.e. $s_{d}^{2}=0,\,s_{b}^{2}=0$), infinitesimal and continuous (dual-)BRST symmetry transformation operators $s_{(d)b}$, and (ii) the discrete duality symmetry transformation operators (13) provide the physical realizations of the Hodge duality * operator of differential geometry.
We now focus on the discrete duality symmetry transformations that exist for the Lagrangian density ${\cal L}_{(\bar{B})}={\cal L}_{(ng)}+{\cal L}_{(fp)}$ [cf. Eqs. (7),(8)]. We observe, in this connection, that most of the fields are common as far as the Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ are concerned. These common basic and auxiliary fields transform, under the discrete duality symmetry transformations for ${\cal L}_{(\bar{B})}$, same as in (13). In the Lagrangian density ${\cal L}_{(\bar{B})}$, there are additional Nakanishi-Lautrup type auxiliary fields (i.e. $\bar{B}_{\mu\nu},\;\bar{\cal B}_{\mu\nu},\;\bar{f}_{\mu},\;F_{\mu}$). These auxiliary fields transform, under the discrete duality symmetry transofmfstions, as:
$\bar{B}_{\mu\nu}\to\pm\,\bar{\cal B}_{\mu\nu},\quad\bar{\cal B}_{\mu\nu}\to\mp\,\bar{B}_{\mu\nu},\quad F_{\mu}\to\pm\,\bar{f}_{\mu},\quad\bar{f}_{\mu}\to\mp\,F_{\mu}.$ (16)
Within the framework of BRST formalism, there is an alternative physical realization of the well-known mathematical relationship: $\delta=\pm\,*\,d\,*$ between the (co-)exterior derivatives $(\delta)d$ of differential geometry in terms of (i) the off-shell nilpotent (i.e. $s_{ab}^{2}=0,\;s_{ad}^{2}=0$), infinitesimal and continuous anti-BRST symmetry transformations $s_{ab}$ [cf. Eq. (10)] and the anti-co-BRST symmetry transformations $s_{ad}$ [cf. Eq. (9)], and (ii) the discrete dualityWe christen the discrete symmetry transformations in (13) and (16) as the duality symmetry transformations because, under these discrete symmetry transformations, we note that (i) the gauge-fixing term of the Abelian 1-form basic gauge field transforms to the kinetic term for the Abelian 3-form basic gauge field and vice-versa, and (ii) the gauge-fixing term for the basic Abelian 3-form gauge field exchanges with the kinetic term for the Abelian 1-form basic gauge field. As stated earlier, the kinetic terms and gauge-fixing terms for the basic Abelian 3-form and 1-form gauge fields owe their mathematical origins to the exterior and dual-exterior derivatives of differential geometry, respectively. symmetry transformations (13) and (16). This statement can be mathematically expressed in the following explicit form, namely;
$s_{ab}\,\Sigma$ $=$ $-\,*\,s_{ad}\,*\,\Sigma,\qquad\Sigma=\beta_{\mu},\;\bar{\beta}_{\mu},\,B_{4},\,B_{5},\qquad s_{ab}\,\widetilde{\Phi}=+\,*\,s_{ad}\,*\,\widetilde{\Phi}$,
$\widetilde{\Phi}$ $=$ $A_{\mu\nu\sigma},\,B_{\mu\nu},\,{\cal B}_{\mu\nu},\,\bar{B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu},\,\bar{C}_{\mu\nu},\,C_{\mu\nu},\,A_{\mu},\,\phi_{\mu},\widetilde{\phi}_{\mu},\bar{f}_{\mu},\,F_{\mu},\,f_{\mu},\,\bar{F}_{\mu},$ (17)
$\bar{C},\,C,\,\bar{C}_{1},\,C_{1},\,\bar{C}_{2},\,C_{2},\,B,\,B_{1},\,B_{2},\,B_{3},$
where the $\mp$ signs in the relationships $s_{ab}\,\Sigma=-\,*\,s_{ad}\,*\,\Sigma$ and $s_{ab}\,\widetilde{\Phi}=+\,*\,s_{ad}\,*\,\widetilde{\Phi}$ are just the opposite of our observations in (15) with the replacement: $\Phi\to\widetilde{\Phi}$. .
In our earlier works [13,14], we have been able to demonstrate that there are total four (anti-)BRST as well as the (anti-)co-BRST invariant CF-type restrictions: $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;{\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$ on our 4D BRST-quntized field-theoretic system. It is very interesting to point out that the following non-trivial anticommutators between the BRST and anti-BRST symmetry transformation operators exist (when they operate on the fields: $A_{\mu\nu\sigma},\,C_{\mu\nu},\,\bar{C}_{\mu\nu}$), namely;
$\{s_{b},\,s_{ab}\}\,A_{\mu\nu\sigma}=\partial_{\mu}(B_{\nu\sigma}+\bar{B}_{\nu\sigma})+\partial_{\nu}(B_{\sigma\mu}+\bar{B}_{\sigma\mu})+\partial_{\sigma}(B_{\mu\nu}+\bar{B}_{\mu\nu}),$
$\{s_{b},\,s_{ab}\}\,C_{\mu\nu}=\partial_{\mu}(f_{\nu}+F_{\nu})-\partial_{\nu}(f_{\mu}+F_{\mu}),$
$\{s_{b},\,s_{ab}\}\,\bar{C}_{\mu\nu}=\partial_{\mu}(\bar{f}_{\nu}+\bar{F}_{\nu})-\partial_{\nu}(\bar{f}_{\mu}+\bar{F}_{\mu}),$ (18)
which turn out to be zero on the subspace of quantum fields where three of the above CF-type restrictions (i.e. $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) are satisfied. It is pertinent to point out, at this juncture, that the absolute anticommutativity (i.e. $\{s_{b},\;s_{ab}=0$) is trivially satisfied for the rest of the fields of the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$. Against the backdrop of our present discussions, it is worthwhile to point out that the following non-trivial anticommutators between the off-shell nilpotent co-BRST and anti-co-BRST symmetry transformation operators $s_{(a)d}$, namely;
$\{s_{d},\,s_{ad}\}\,A_{\mu}=+\,\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}\big({\cal B}^{\sigma\rho}+\bar{\cal B}^{\sigma\rho}\big),$
$\{s_{d},\,s_{ad}\}\,C_{\mu\nu}=+\,\big[\partial_{\mu}(f_{\nu}+F_{\nu})-\partial_{\nu}(f_{\mu}+F_{\mu})\big],$
$\{s_{b},\,s_{ad}\}\,\bar{C}_{\mu\nu}=+\,\big[\partial_{\mu}(\bar{f}_{\nu}+\bar{F}_{\nu})-\partial_{\nu}(\bar{f}_{\mu}+\bar{F}_{\mu})\big],$ (19)
turn out to be zero provided we use the three specific CF-type restrictions (i.e. ${\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) which define the subspace of qauntum fields where the co-BRST and anti-co-BRST symmetry transformation operators absolutely anticommute (i.e. $\{s_{d},\;s_{ad}\}=0$). The absolute anticommutativity (i.e. $\{s_{d},\;s_{ad}\}=0$) is satisfied trivially for the rest the quantum fields of our 4D BRST-quantized theory. Hence, we finally conclude that the absolute anticommutativity (i.e. $\{s_{b},\;s_{ab}\}=0$ and $\{s_{d},\;s_{ad}\}=0$) between the off-shell nilpotent (i.e. $s_{(a)b}^{2}=0,\,s_{(a)d}^{2}=0$) (anti-)BRST ($s_{(a)b}$) and (anti-)dual-BRST ($s_{(a)d}$) symmetry transformation operators are satisfied provided we utilize the specific set of only three CF-type restrictions (out of the total four).
We end this section with a few key concluding remarks. First of all, we observe that the following algebraic operator structures (in the terminology of the anti-commutators)
$\big\{s_{b},\;s_{ad}\big\}=0,\qquad\big\{s_{d},\;s_{ab}\big\}=0,$ (20)
are satisfied modulo a set of U(1) vector gauge symmetry-type transformations. To be precise, we observe that the following are true, namely;
$\big\{s_{b},\;s_{ad}\big\}\bar{\beta}_{\mu}=\partial_{\mu}\big(B_{2}-B_{3}\big),\quad\big\{s_{b},\;s_{ad}\big\}\phi_{\mu}=-\,\partial_{\mu}B_{4},\quad\big\{s_{b},\;s_{ad}\big\}\widetilde{\phi}_{\mu}=-\,\partial_{\mu}B_{4},$
$\big\{s_{d},\;s_{ab}\big\}\beta_{\mu}=\partial_{\mu}\big(B_{2}+B_{3}\big),\quad\big\{s_{d},\;s_{ab}\big\}\phi_{\mu}=+\,\partial_{\mu}B_{5},\quad\big\{s_{d},\;s_{ab}\big\}\widetilde{\phi}_{\mu}=-\,\partial_{\mu}B_{5}.$ (21)
Furthermore, we would like to add that the anticommutators in (20) do not define a new set of bosonic symmetry transformations because the Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ do not (i) remain invariant under (20), and/or (ii) transform to the total spacetime derivatives under (20). It is worthwhile to mention that the anticommutators (20) are true for the rest of the fields of our theory [except the ones that are present in (21)]. It is very interesting to point out, in our present context, that if we make the choice: $B_{2}=B_{3}=B_{4}=B_{5}=0$, we find that the anticommutators in (20) exactly become zero. However, even under this limit, the off-shell nilpotent symmetry transformations $s_{(a)b}$ and $s_{(a)d}$ as well as the bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ exist. Second, the following non-trivial anticmmmutators between the off-shell nilpotent symmetry operators, namely;
$\big\{s_{b},\;s_{d}\big\}=s_{\omega},\;\;\qquad\;\;\big\{s_{ad},\;s_{ab}\big\}=s_{\bar{\omega}},$ (22)
define the two non-trivial bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$, respectively (for our 4D BRST-quantized field-theoretic system). Third, we would like to point out that some of the key relationships between the co-BRST and BRST symmetry transformations have not been captured in equation (14). Taking into account the total discrete duality symmetry transformations (13) and (16), it is straightforward to check that:
$s_{d}\bar{f}_{\mu}=-\,*\,s_{b}\,*\bar{f}_{\mu}=\partial_{\mu}B_{5},\qquad s_{d}F_{\mu}=-\,*\,s_{b}\,*\,F_{\mu}=-\,\partial_{\mu}B_{3},$
$s_{d}\bar{\cal B}_{\mu\nu}=-\,*\,s_{b}\,*\,\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\bar{F}_{\nu}-\partial_{\nu}\bar{F}_{\mu}).$ (23)
Finally, it is the peculiarity of our 4D BRST-quantized theory that the reciprocal relationships of (14) and (17) also exist and they are found to be true. In other words, it can be verified clearly that (i) the relationships $s_{b}\,\Sigma=+\,*\,s_{d}\,*\,\Sigma$ and $s_{b}\,\Phi=-\,*\,s_{d}\,*\,\Phi$, and (ii) the relationships $s_{ad}\,\Sigma=-\,*\,s_{ab}\,*\,\Sigma$ and $s_{ad}\,\Phi=+\,*\,s_{ab}\,*\,\Phi,$ are also true. However, the actual mapping between the nilpotent symmetry and cohomological operators $(\delta)d$ would be dictated by the ghost number considerations (cf. our section five for more discussions).
4 Bosonic Symmetry Transformations: Uniqueness Property and the Full Set of CF-Type Restrictions
The non-nilpotent (i.e. $s_{\omega}^{2}\neq 0$) bosonic symmetry operator $s_{\omega}$ [cf. Eq. (22)] can operate on the individual fields of this Lagrangian density ${\cal L}_{(B)}$. The following infinitesimal, continuous and non-nilpotent symmetry transformations under $s_{\omega}=\{s_{b},\;s_{d}\}$, namely;
$s_{\omega}A_{\mu\nu\sigma}=\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\rho}B+\Big(\partial_{\mu}\,{\cal B}_{\nu\sigma}+\partial_{\nu}\,{\cal B}_{\sigma\mu}+\partial_{\sigma}\,{\cal B}_{\mu\nu}\Big),$
$s_{\omega}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}B^{\sigma\rho}+\partial_{\mu}B_{1},\qquad s_{\omega}\bar{\beta}_{\mu}=+\,\partial_{\mu}B_{5},\qquad s_{\omega}\beta_{\mu}=+\,\partial_{\mu}B_{4},$
$s_{\omega}C_{\mu\nu}=-\,\big(\partial_{\mu}f_{\nu}-\partial_{\nu}f_{\mu}\big),\qquad s_{\omega}\bar{C}_{\mu\nu}=+\,\big(\partial_{\mu}\bar{F}_{\nu}-\partial_{\nu}\bar{F}_{\mu}\big),$
$s_{\omega}\Big[B,\,B_{1},B_{2},B_{3},B_{4},B_{5},\,\phi_{\mu},\,\widetilde{\phi}_{\mu},\,f_{\mu},\,\bar{F}_{\mu},\,C,\bar{C},\,C_{1},\bar{C}_{1},C_{2},\bar{C}_{2},B_{\mu\nu},\,{\cal B}_{\mu\nu}\Big]=0,$ (24)
lead to the transformation of the Lagrangian density ${\cal L}_{(B)}$ as follows
$s_{\omega}\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,B_{\nu\sigma}$
$-\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,{\cal B}_{\nu\sigma}+B_{4}\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}B_{4}$
$+\,\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big)\,\bar{F}_{\nu}+\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,f_{\nu}\Big]+\partial_{\mu}\Big[\big(B_{1}\,\partial^{\mu}B-B\,\partial^{\mu}B_{1}\big)\Big],$ (25)
thereby rendering the action integral $S=\int d^{4}x\,{\cal L}_{(B)}$ invariant (i.e. $s_{\omega}S=0$). The transformation in (25) can be easily derived by using our observations in equations (5) and (6). In other words, we have $s_{\omega}{\cal L}_{(B)}=\big(s_{b}\,s_{d}+s_{d}\,s_{b}\big)\,{\cal L}_{(B)}$ which implies the transformation (25)
Against the backdrop of the above discussions, we note that the operation of the bosonic symmetry transformation operator $s_{\bar{\omega}}=\{s_{ab},\;s_{ad}\}$ on the individual fields of the Lagrangian density ${\cal L}_{(\bar{B})}$ leads to the following:
$s_{\bar{\omega}}A_{\mu\nu\sigma}=-\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\rho}B+\Big(\partial_{\mu}\,\bar{\cal B}_{\nu\sigma}+\partial_{\nu}\,\bar{\cal B}_{\sigma\mu}+\partial_{\sigma}\,\bar{\cal B}_{\mu\nu}\Big),$
$s_{\bar{\omega}}A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}\bar{B}^{\sigma\rho}-\partial_{\mu}B_{1},\qquad s_{\bar{\omega}}\bar{\beta}_{\mu}=-\,\partial_{\mu}B_{5},\qquad s_{\bar{\omega}}\beta_{\mu}=-\,\partial_{\mu}B_{4},$
$s_{\bar{\omega}}C_{\mu\nu}=-\,\big(\partial_{\mu}F_{\nu}-\partial_{\nu}F_{\mu}\big),\qquad s_{\bar{\omega}}\bar{C}_{\mu\nu}=+\,\big(\partial_{\mu}\bar{f}_{\nu}-\partial_{\nu}\bar{f}_{\mu}\big),$
$s_{\bar{\omega}}\Big[B,\,B_{1},B_{2},B_{3},B_{4},B_{5},\,\phi_{\mu},\,\widetilde{\phi}_{\mu},\,\bar{f}_{\mu},\,F_{\mu},\,C,\bar{C},\,C_{1},\bar{C}_{1},C_{2},\bar{C}_{2},\bar{B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu}\Big]=0.$ (26)
Taking into account the sum of transformations (24) and (26), we obtain the following:
$\big(s_{\omega}+s_{\bar{\omega}}\big)A_{\mu\nu\sigma}=+\,\Big[\partial_{\mu}\,\big({\cal B}_{\nu\sigma}+\bar{\cal B}_{\nu\sigma}\big)+\partial_{\nu}\,\big({\cal B}_{\sigma\mu}+\bar{\cal B}_{\sigma\mu}\big)+\partial_{\sigma}\,\big({\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}\big)\Big],$
$\big(s_{\omega}+s_{\bar{\omega}}\big)A_{\mu}=\dfrac{1}{2}\,\varepsilon_{\mu\nu\sigma\rho}\,\partial^{\nu}\big(B^{\sigma\rho}+\bar{B}^{\sigma\rho}\big),$
$\big(s_{\omega}+s_{\bar{\omega}}\big)C_{\mu\nu}=-\,\Big[\big(\partial_{\mu}(f_{\nu}+F_{\nu})-\partial_{\nu}(f_{\mu}+F_{\mu})\big)\Big],$
$\big(s_{\omega}+s_{\bar{\omega}}\big)\bar{C}_{\mu\nu}=+\,\Big[\big(\partial_{\mu}(\bar{f}_{\nu}+\bar{F}_{\nu})-\partial_{\nu}(\bar{f}_{\mu}+\bar{F}_{\mu})\big)\Big],$
$\big(s_{\omega}+s_{\bar{\omega}}\big)\Big[B,\,B_{1},B_{2},B_{3},B_{4},B_{5},\phi_{\mu},\widetilde{\phi}_{\mu},f_{\mu},\bar{F}_{\mu},\bar{f}_{\mu},F_{\mu},\beta_{\mu},\bar{\beta}_{\mu},$
$C,\bar{C},\,C_{1},\bar{C}_{1},C_{2},\bar{C}_{2},B_{\mu\nu},\,{\cal B}_{\mu\nu},\bar{B}_{\mu\nu},\,\bar{\cal B}_{\mu\nu}\Big]=0.$ (27)
It is crystal clear that the relationship $s_{\omega}\,+\,s_{\bar{\omega}}=0$ is true provided we consider our whole BRST-quantized theory on the space of quantum fields where all the four CF-type restrictions (i.e. $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;{\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) are satisfied together. This observation should be contrasted against the backdrop of our observations in the context of the proof of the absolute anticommutativity (i) between the off-shell nilpotent BRST and anti-BRST symmetry transformation operators [cf. Eq. (18)], and (ii) between the off-shell nilpotent co-BRST and anti-co-BRST symmetry transformation operators [cf. Eq. (19)], where only three (out of the total four) specific types of CF-type restrictions have been invoked. We conclude, ultimately, that there is a unique bosonic symmetry transformation operator (e.g. $s_{\omega}$) in our theory.
As far as the invariance of the action integral $S=\int d^{4}x\,{\cal L}_{(\bar{B})}$ (corresponding to the Lagrangian density ${\cal L}_{(\bar{B})}$) w.r.t the bosonic symmetry transformation operator $s_{\bar{\omega}}$ is concerned, we observe that ${\cal L}_{(\bar{B})}$ transforms to the following total spacetime derivative, namely;
$s_{\bar{\omega}}\,{\cal L}_{(\bar{B})}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,\bar{B}^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,\bar{\cal B}_{\nu\sigma}$
$-\,\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}-B_{4}\,\partial^{\mu}B_{5}+B_{5}\,\partial^{\mu}B_{4}$
$-\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,F_{\nu}-\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\bar{f}_{\nu}\Big]+\partial_{\mu}\Big[\big(B\,\partial^{\mu}B_{1}-B_{1}\,\partial^{\mu}B\big)\Big].$ (28)
In addition to our bosonic symmetry observations in (25) and (28), we note that the following bosonic transformations are true, namely;
$s_{\omega}\,{\cal L}_{(\bar{B})}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,\bar{\cal B}_{\nu\sigma}$
$-\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}+B_{4}\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}B_{4}$
$-\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,\bar{f}_{\nu}-\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,F_{\nu}\Big]+\partial_{\mu}\Big[\big(B_{1}\,\partial^{\mu}B-B\,\partial^{\mu}B_{1}\big)\Big]$
$+\;\partial_{\mu}\big[B_{\nu\sigma}+\bar{B}_{\nu\sigma}-(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu})\big]\,\big(\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big)$
$-\;\partial_{\mu}\big[\bar{f}_{\nu}+\bar{F}_{\nu}-\partial_{\nu}\bar{C}_{1}\big]\,\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big),$ (29)
$s_{\bar{\omega}}\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,B_{\nu\sigma}$
$-\,\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,{\cal B}_{\nu\sigma}-B_{4}\,\partial^{\mu}B_{5}+B_{5}\,\partial^{\mu}B_{4}$
$+\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,f_{\nu}+\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\bar{F}_{\nu}\Big]+\partial_{\mu}\Big[\big(B\,\partial^{\mu}B_{1}-B_{1}\,\partial^{\mu}B\big)\Big],$
$-\;\partial_{\mu}\big[{\cal B}_{\nu\sigma}+\bar{\cal B}_{\nu\sigma}-(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu})\big]\,\big(\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big)$
$-\;\partial_{\mu}\big[f_{\nu}+F_{\nu}-\partial_{\nu}C_{1}\big]\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big),$ (30)
which establish that the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ respect both the bosonic symmetry transformations $s_{\omega}$ and $s_{\bar{\omega}}$ provided our whole theory is considered on the space of quantum fields where all the four CF-type restrictions are satisfied together. Furthermore, against the backdrop of our observations in (27), we note the following interesting observations for the transformations of the Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$, namely;
$\big(s_{\omega}+s_{\bar{\omega}}\big)\,{\cal L}_{(\bar{B})}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big\}\,\bar{\cal B}_{\nu\sigma}$
$-\big\{\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big\}\,\bar{B}_{\nu\sigma}+B_{4}\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}B_{4}$
$-\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,\bar{f}_{\nu}-\big(\partial^{\mu}\bar{F}^{\nu}-\partial^{\nu}\bar{F}^{\mu}\big)\,F_{\nu}\Big]+\partial_{\mu}\Big[\big(B_{1}\,\partial^{\mu}B-B\,\partial^{\mu}B_{1}\big)\Big]$
$+\;\partial_{\mu}\Big[B_{\nu\sigma}+\bar{B}_{\nu\sigma}-(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu})\Big]\,\big(\partial^{\mu}\,{\cal B}^{\nu\sigma}+\partial^{\nu}\,{\cal B}^{\sigma\mu}+\partial^{\sigma}\,{\cal B}^{\mu\nu}\big)$
$-\;\partial_{\mu}\Big[\bar{f}_{\nu}+\bar{F}_{\nu}-\partial_{\nu}\bar{C}_{1}\Big]\,\big(\partial^{\mu}f^{\nu}-\partial^{\nu}f^{\mu}\big),$ (31)
$\big(s_{\omega}+s_{\bar{\omega}}\big)\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\,\Big[\big\{\partial^{\mu}\,\bar{\cal B}^{\nu\sigma}+\partial^{\nu}\,\bar{\cal B}^{\sigma\mu}+\partial^{\sigma}\,\bar{\cal B}^{\mu\nu}\big\}\,B_{\nu\sigma}$
$-\,\big\{\partial^{\mu}\,\bar{B}^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,\bar{B}^{\mu\nu}\big\}\,{\cal B}_{\nu\sigma}-B_{4}\,\partial^{\mu}B_{5}+B_{5}\,\partial^{\mu}B_{4}$
$+\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big)\,f_{\nu}+\big(\partial^{\mu}F^{\nu}-\partial^{\nu}F^{\mu}\big)\,\bar{F}_{\nu}\Big]+\partial_{\mu}\Big[\big(B\,\partial^{\mu}B_{1}-B_{1}\,\partial^{\mu}B\big)\Big],$
$-\;\partial_{\mu}\Big[{\cal B}_{\nu\sigma}+\bar{\cal B}_{\nu\sigma}-(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu})\Big]]\,\big(\partial^{\mu}\,B^{\nu\sigma}+\partial^{\nu}\,B^{\sigma\mu}+\partial^{\sigma}\,B^{\mu\nu}\big)$
$-\;\partial_{\mu}\Big[f_{\nu}+F_{\nu}-\partial_{\nu}C_{1}\Big]]\,\big(\partial^{\mu}\bar{f}^{\nu}-\partial^{\nu}\bar{f}^{\mu}\big).$ (32)
Thus, ultimately, we have been able to establish the uniqueness of the bosonic symmetry transformation operator at the level of symmetry transformations on (i) the individual fields [cf. Eq. (27)], and (ii) the coupled (but equivalent) Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ [cf. Eqs. (31),(32)]. We would like add here that this key observation is true only on the space of quantum fields where all the four CF-type restrictions are satisfied. It is interesting to end this section with the final remark that the unique (i.e. $s_{\omega}=-\,s_{\bar{\omega}}$) bosonic symmetry transformation operator commutes (i.e. $s_{\omega},\;s_{(a)b}]=0,\;[s_{\omega},\;s_{(a)d}]=0$) with all the four nilpotent symmetry transformation operators (i.e. $s_{(a)b},\;s_{(a)d}$) of our theory.
5 Conclusions
In our present endeavor, we have very briefly sketched the existence of the infinitesimal, continuous and off-shell nilpotent (i.e. $s_{(a)b}^{2}=0,s_{(a)d}^{2}=0$) versions of the (anti-)BRST and (anti-)co-BRST symmetry transformation operators (i.e. $s_{(a)b}$ and $s_{(a)d}$, respectively) for the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ that describe the 4D BRST-quantized combined field-theoretic system of the free Abelian 3-form and 1-form gauge theories. In addition, we have given a great deal of importance to the existence of a couple of discrete duality symmetry transformations (in the context of our 4D BRST-quantized theory) and have highlighted their roles in the algebraic structures that are obeyed by the four off-shell nilpotent, infinitesimal and continuous symmetry transformation operators that exist for the coupled (but equivalent) Lagrangian densities of our theory. To be precise, we have provided the physical realizations [cf. Eqs. (14),(17)] of the well-known mathematical relationship $\delta=\pm\,*\,d\,*$ between the nilpotent (i.e. $d^{2}=0,\delta^{2}=0$) (co-)exterior derivatives $(\delta)d$ of differential geometry in terms of the interplay between (i) the discrete duality symmetry transformation operators [cf. Eq. (13)], and (ii) the infinitesimal, continuous and off-shell nilpotent symmetry transformation operators [cf. Eqs. (3),(4),(9),(10)].
Out of the total six anticommutators (that can be constructed from the four nilpotent symmetry operators), we have shown that (i) the two specific anticommutators (i.e. $\{s_{d},\;s_{ab}\}=0$ and $\{s_{ad},\;s_{b}\}=0$) are trivially zero on their own for all the quantum fields of our 4D BRST-quantized theory [except for the fields $\beta_{\mu},\;\bar{\beta}_{\mu},\;\phi_{\mu}$ and $\widetilde{\phi}_{\mu}$ that are present in equation (21) which transform up to a set of U(1) vector gauge symmetry-type transformations], (ii) another two anticommutators (i.e. $\{s_{d},\;s_{ad}\}=0$ and $\{s_{ab},\;s_{b}\}=0$) are zero provided we use specific set of three (out of the total four) CF-type restrictions on our theory, and (iii) the rest of the two anticommutators (i.e. $\{s_{d},\;s_{b}\}=s_{\omega}$ and $\{s_{ab},\;s_{ad}\}=s_{\bar{\omega}}$) define a set of two bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$. We have established, in our present investigation, that there exists a unique bosonic symmetry operator (i.e. $s_{\omega}=-\,s_{\bar{\omega}}$) on the subspace (of the total field space) of quantum fields which is precisely defined by all the four CF-type restrictions (i.e. $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;{\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) that exist on our 4D BRST-quantized field-theoretic system [cf. Eqs. (27),(31),(32)].
We find that (i) the algebraic structures of the symmetry transformation operators (in our section three and four), and (ii) the ghost number considerations (in the context of off-shell nilpotent symmetry operators) are such that we obtain a two-to-one mapping [i.e. $(s_{b},\;s_{ad})\Rightarrow d,\;(s_{d},\;s_{ab})\Rightarrow\delta,\;(s_{\omega},\;s_{\bar{\omega}})\Rightarrow\Delta$] between (i) the symmetry transformation operators, and (ii) the de Rham cohomological operators of differential geometry (where the discrete duality symmetry operators [cf. Eq. (13)] provide the physical realization(s) of the Hodge duality * operator). The study of the BRST-quantized field-theoretic systems has been mathematically and physically useful in the sense that we have been able to establish that the 2D (non-)Abelian gauge theories (without any interaction with the matter fields) provide [17] a new type of topological field theory (TFT) that captures a few aspects of the Witten-type TFT [18] and some salient features of the Schwarz-type TFT [19]. In addition, such studies have led to the existence of “exotic” fields with the negative kinetic terms (see, e.g. [15] and references therein) that are a set of possible candidates for (i) the “phantom” and/or “ghost” fields of the cyclic, bouncing and self-accelerated cosmological models of the Universe (see, e.g. [20-22]), and (ii) the dark energy/dark matter (see, e.g. [23,24]).
In our future endeavor, we plan to derive (i) all the Noether conserved charges corresponding to the continuous symmetry transformations of our 4D BRST-quantized theory, and (ii) the algebraic structures satisfied by the appropriate conserved charges and study the mapping between the appropriate conserved charges and the de Rham cosmological operators of differential geometry. In other words, we wish to prove that (i) our present 4D BRST-quantized field-theoretic system, and (ii) the Stückelberg-modified massive version of our present system, are a set of the tractable examples for Hodge theory.
Data Availability Statement
No new data were created or analyzed in this study.